9x^3-3x^2/3x
I tried doing this but its hard for me to understand the simplifying. All I got was 3(3x^3-x^2)/3x can u help me?

Answer by richwmiller(17219)

I will assume that everything is over 3x
Look again for something else to pull out of 3(3x^3-x^2)
Remember that x^3=x*x*x and x^2=x*x
The denominator 3x should be a hint
after you pull out one more item then that has another set of factors.
Hint :Can you factor (6w^2-w)? yes of course you can.